OpenStudy (anonymous):

Wheel Industries is considering a three-year expansion project. The project requires an initial investment of $1.5 million. The project will use straight-line depreciation method. The project has no salvage value. It is estimated that the project will generate additional revenues of $1.2 million per year and has annual costs of $600,000. The tax rate is 35 percent. Calculate the cash flows for the project. If the discount rate were 6 percent, calculate the NPV of the project, then offer your thoughts as to whether this is an economically acceptable project to undertake. Clinton Co. has jus

OpenStudy (anonymous):

After tax cash flow per year = (1,200,000 - 600,000) * (1-0.35) = 390,000 Value of the project = (390,000) / 0.06 = $6,500,000 NPV = 6,500,000 - 1,500,000 = $5,000,000

OpenStudy (anonymous):

The npv would be 390/1.06 + 390/1.06^2 + 390/1.06^3